OpenStudy (anonymous):

integral of (1)/(xsqrt(4x^2-9))

OpenStudy (anonymous):

... this is too hard 4 m3 what ish dis? so hard!!

OpenStudy (campbell_st):

use substituttion let \[u = 4x^2 - 9\] and then \[\frac{du}{dx} = 8x \] so \[\frac{1}{8} du = x ~ dx\] so the substitution becomes \[\int\limits \frac{1}{8} u^{-\frac{1}{2}} du~~or~~\frac{1}{8} \int\limits u^{-\frac{1}{2}} ~du\] hope it helps

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

I'm actually suppose to use trig substitution on this problem. That's why I was confused